Red drives 20 miles south on Route 295 towards Warwick at a rate of 60 miles per hour and it takes 20 minutes. Rusty takes the same trip during a snowstorm at a rate of 30 miles per hour. What is his time?
Answer #1

Speed is defined as the rate of change of distance or the distance covered per unit time .

Thus formula of speed becomes :

Speed = Distance/Time ........1

The distance travelled by Red is 20 miles. He drove at the speed of 60 miles per hour and thus time taken by him to cover 20 miles at the speed of 60 miles per hour is given to be 20 minutes.

Now Rusty travels the same distance but at a different speed. His speed is 30 miles per hour. We need to calculate the time taken by Rusty to cover a distance of 20 miles at a speed of 30 miles per hour .

Since Speed = Distance/Time ,

So Time = Distance/Speed

Thus time taken by Rusty = 20 miles/ 30 miles per hour = (20/30) hours = (2/3) hours

As 1 hour = 60 minutes so his time in minutes would be :

Time taken by Rusty = (2/3) hours = (2/3) × 60 minutes = 40 minutes